A very long uniform line of charge has charge per unit length 4.80 µC/m and lies along the x￾axis. A second long uniform line of charge has charge per unit length – 2.40 µC/m and is parallel


to the -axis at y = 0.400m. What is the net electric field (magnitude and direction) at the following


points on the y-axis: (a) y = 0.200 m and (b) y = 0.600 m?



1
Expert's answer
2022-04-17T17:06:42-0400

a)

"E_1=2k(\frac {\tau_1}{0.2}+\frac{\tau_2}{0.2})=0,"

b)

"E_2=2k(\frac {\tau_1}{0.6}+\frac{\tau_2}{0.2})=-72~\frac{kN}C."